spectral problem its all one problem
1) identify the compounds for each group of spectra (mass, proton, carbon and IR)
2) assign the proton and carbon and IR spectra. identify at least two possible fragmentation on the mass spectrum
#8 1H 5.5 5.0 20 ag 200 180 160 140 120 100 80 0 40 20
C-69.8: H-11.6 -18.6 100 - MS-W-11 80 60 40- 20- 10 20 30 40 50 60 70 80 Rcfon c-cac-c-R R7C=0

the values is whats there from the 3rd graph thats all i was given
Answer #1

MS data:

M+ (parent peak) = 86

Percent of carbon = 69.8%

Percent of hydrogen = 11.6%

Percent of oxygen = 18.6%

To calculate empirical formula

Converting mass of each element to moles using molar mass

molar mass of carbon = 69.8/12 = 5.8

molar mass of hydrogen = 11.6 / 1 = 11.6

molar mass of oxygen = 18.6 / 16 = 1.16

Dividing each mole value by the smallest number of moles calculated. Round to the nearest whole number.

Number of carbon = 5.8/1.16 = 5

Number of hydrogen = 11.6/1.16 = 10

Number of oxygen = 1.16/1.16 = 1

Empirical formula = C5H10O

Molecular weight of empirical formula is 86 and hence molecular formula is same that is C5H10O

Degree of Unsaturation (DOU) is also known as Double Bond Equivalent. If the molecular formula is given, DoU can be calculated as follows

DoU   =   2C+2+N−X−H

               2

C is the number of carbons

N is the number of nitrogens

X is the number of halogens (F, Cl, Br, I)

H is the number of hydrogens

Molecular formula = C5H10O

Sites of unsaturation = 1

1H NMR data:

NMR peak value (ppm)

Protons

Indication

5.4

1H

proton attached to CH2

4.1

2H

CH2 attached to electronegative atom such as oxygen

2.4

1H

Broad peak due to exchangeable proton (OH)

1.8

3H

CH3 attached to quaternary carbon

1.7

3H

CH3 attached to quaternary carbon

13C NMR data:

NMR peak value (ppm)

Indication

110

Alkene carbon

105

Alkene carbon attached to proton (CH)

64

Aliphatic CH2 attached to oxygen

42

Aliphatic CH3

40

Aliphatic CH3

IR data:

IR peak (cm-1)

Indication

3200 broad peak

O-H stretch

2900

Aliphatic C-H stretch

1650

C=C stretch of alkene

All above data confirms following structure

-OH 3-methylbut-2-en-1-ol Exact Mass: 86.07